### v1

# Systematic study of confinement induced effects on atomic electronic structure

Abstract
We point out that although a litany of studies have been published on atoms in hard-wall confinement, they have either not been systematic, having only looked at select atoms and/or select electronic configurations, or they have not used robust numerical methods. To remedy the situation, we perform in this work a methodical study of atoms in hard-wall confinement with the HelFEM program, which employs the finite element method (FEM) that trivially implements the hard-wall potential, guarantees variational results, and allows easily finding the numerically exact solution.
We systematically examine low lying configurations of the H–Xe atoms and their monocations, and investigate how the states—especially the ground state—behave as a function of the position of the hard-wall boundary. Our fully numerical calculations are based on non-relativistic density functional theory (DFT), employing the Perdew–Burke–Ernzerhof (PBE) exchange-correlation functional and spherically averaged densities. We consider calculations with both spin-polarized as well as spin-restricted densities, and demonstrate that spin-polarization effects are significant in open shell configurations, even though some previous studies have only considered the spin-restricted model.
We demonstrate the importance of considering ground state changes for confined atoms by computing the ionization radii for the H–Xe atoms and observe significant differences to earlier studies. Confirming previous observations, we identify electron shifts on the outermost shells for a majority of the elements: valence $s$ electrons are highly unfavored under strong confinement, and the high-lying $3d$ and $4f$ orbitals become occupied in atoms of periods 2–3 and 4, respectively.
1 Introduction
The pioneering study of Michels et al.1 was concerned on the question of how atomic polarizabilities—which have an important role in chemical bonding—evolve as a function of pressure. Published in 1937, the study examined the hydrogen atom enclosed in an impenetrable sphere of a variable radius to simulate pressure effects. Such frugal models offer a great tool for understanding basic physical effects precisely due to their simplicity: Michels et al.1 found that hydrogen becomes less polarizable in increasing pressure.
An interest in analogous studies of atoms in hard-wall confinement continues to this day. A multitude of studies have been published on the very first atoms of the periodic table; see ref. 2 for a review dedicated to the H and He atoms confined by a finite or infinite spherical barrier with 200 references. Yet, few studies have considered many-electron atoms, or reported systematic calculations on the periodic table.
In hard-wall confinement, the wave function must vanish beyond the hard-wall boundary, and this can only be accomplished within a basis set with finite support. Although studies on confined atoms employing basis sets of an analytic form—such as Gaussian or Slater type orbital basis sets—have been published, they have finite tails in the forbidden region. Truncated Gaussian or Slater orbital basis sets do not have this issue, but they also suffer from basis set truncation errors. Our focus here is on fully numerical studies that are able to both accurately describe confinement effects as well as recover the numerically exact wave function.
We will now briefly review the literature we found that fulfills our key criteria at least in part: studies mentioned herein have examined atoms with more than 4 electrons, and employed a numerically robust method.
Boeyens3 performed fully numerical Hartree–Fock calculations to determine the ionization radii of compressed atoms. However, Boeyens3 did not employ a proper hard-wall potential. To approximate a hard-wall boundary at $r=r_{\mathrm{c}}$, Boeyens3 multiplied the wave function by the smooth step function $\Theta(r)=\exp\left[-\left(r/r_{c}\right)^{20}\right]$ at every self-consistent field (SCF) iteration, before the normalization of the updated wave function was carried out. Boeyens3 employed standard electronic configurations for the elements, and thus did not consider changes in the occupations as a function of the confinement.
[…]
Sarkar et al.6 studied the effects of confinement on the chemical reactivity of the nitrogen atom. Sarkar et al.6 compared the method of Boeyens3 to the true hard-wall potential, and found the arising chemical reactivities to yield the same qualitative trends. Moreover, Sarkar et al.6 found that different methods to calculate the reactivity yield dissimilar results, and that different electronic occupations also lead to different chemical reactivities.
Connerade et al.7 studied the filling of shells in the $3d$ and $4d$ elements as well as their cations in average-of-configurations Hartree–Fock calculations in a spherical cavity surrounded by a $10E_{h}$ potential wall. They found that orbital filling becomes more regular for successive rows with increasing pressure. In specific, the $s$$d$ competition disappears for these elements, as the $(n-1)d$ orbitals become energetically more favourable than the $ns$ orbitals under confinement. Connerade and coworkers have also reported studies on some individual atoms under confinement. For instance, Connerade and Dolmatov examined the Cr atom, while Connerade and Semaoune performed relativistic calculations on the La and Cs atoms in refs. 9 and 10, respectively.
A number of studies of atoms in hard-wall confinement have been carried out at the DFT level by Garza et al.. Garza et al.11 reported a finite differences method (FDM) implementation for solving the atomic DFT equations in the presence of a hard-wall boundary, and studied the Ne and Na (and He) atoms in exchange-only calculations. In a follow-up study, Garza et al.12 examined the electrostatic potential of the unconfined and hard-wall confined Li, Na, K, and Rb atoms with DFT in order to study changes in the atoms’ shell structure, and found that shell structure is gradually lost in increasing confinement. Similarly, Sen et al.13 determined atomic ionization radii for the He–Ca atoms with hard-wall DFT calculations. As none of the studies by Garza et al.11, Garza et al.12, or Sen et al.13 discuss occupations, we can only assume that the ground state electron configuration of the unconfined atom was used in their calculations. In contrast, the study on the chemical reactivity indices of the compressed Li, Na, and K; Be, Mg, and Ca; N, P, and As; and Ne, Ar, and Kr atoms by Garza et al.14, which was likewise carried out with hard-wall DFT, explicitly mentions finding the lowest electronic configuration for each confinement radius.
Continuing with the Garza group, Guerra et al.15 modeled pressure effects on the electronic properties of Ca, Sr, and Ba with hard-wall DFT, and found that their ground state changes from the $ns^{2}$ configuration to the $ns^{1}(n-1)d^{1}$ and $ns^{0}(n-1)d^{2}$ configurations in increasing pressure. Confinement effects on the spin potential of first-row transition metal cations were investigated at the hard-wall DFT level of theory by Lozano-Espinosa et al.16.
Pašteka et al.17 studied the C and K (and He) atoms in hard confinement with the finite element method (FEM) at the complete active space self-consistent field18, 19 (CASSCF) level of theory, which represent the most sophisticated calculations in the studies we found. Also Pašteka et al.17 found changes in atomic ground state configurations under confinement: the ground state configuration of carbon changes from $1s^{2}2s^{2}2p^{2}$ in the free atom to $1s^{2}2p^{4}$ in strong confinement.
As this literature overview demonstrates, although several fully numerical studies have been published on atoms in confinement, they have mostly focused on small groups of similar atoms. The study of Boeyens3 was the most complete, considering non-relativistic calculations on the first 105 atoms of the periodic table, but the study relied on an ad hoc implementation of confinement, and—like in many other studies mentioned above—changes in the ground state as a function of confinement were not considered. Studies that considered changes in the ground state are likewise limited: Connerade et al.7 only considered $3d$ and $4d$ elements, Garza et al.14 studied only atoms of groups 1, 2, 15, and 18, and Guerra et al.15 only studied heavy atoms of group 2.
The only study that appears to have systematically investigated effects of confinement on the electronic structure of atoms so far is the study of Rahm et al.20, which employed the extreme pressure polarizable continuum model21 (XP-PCM) to study the first 93 atoms under pressure, employing Gaussian basis sets with scalar relativistic corrections and the Perdew–Burke–Ernzerhof (PBE) functional,22, 23 as well as its hybrid with 25% exact exchange, the PBE0 functional.24, 25 Rahm et al.20 identified electron shifts, $s\to p$, $s\to d$, $s\to f$, and $d\to f$, as essential chemical and physical consequences of compression.
As Gaussian basis sets cannot be expected to deliver reliable results in strong confinement,17 there is clearly a hole in the literature regarding systematic fully numerical studies covering (a significant fraction of) the periodic table.
As is clear from the above discussion, changes in the electronic structure of atoms are an expected effect of confinement, and they thus need to be taken into account by considering changes in the electronic occupations as a function of confinement. In this work, we will address this major shortcoming of the existing literature with a systematic study on atoms in hard-wall confinement within a fully numerical approach, while also thoroughly considering the effect of the confinement on the ground state electron configuration.
Our results confirm previous partial findings of the fully numerical studies of Connerade et al.7, Guerra et al.15, and Pašteka et al.17 as well as the thorough Gaussian basis study of Rahm et al.20: confinement induced electron shifts can be observed for a majority of the elements. It is important to note here that all of these studies have been carried out with different methodologies. This suggests that the findings of this work are indeed reliable despite the arguably simplistic level of theory employed herein.
The layout of this work is the following. The form of the Hamiltonian including the hard-wall potential, as well as the level of theory is briefly discussed in section 2. The computational details including the employed numerical approach are presented in section 3. The results are presented in section 4. We first demonstrate that our method works by computing ionization energies and radii in section 4.1 and section 4.2 respectively. Then, we perform a systematic study of the behavior of various low lying configurations of the H–Xe atoms in hard-wall confinement. The full set of raw data as well as a detailed analysis of the results is included in the Supporting Information (SI); in section 4.3 we present the highlights from this analysis. The article concludes in a summary and discussion in section 5. Atomic units are employed throughout the work.
2 Theory
The Hamiltonian for the confined atoms reads
$\hat{H}=\hat{H}_{0}+V_{c}(r),$ (1)
where
$\hat{H}_{0}=-\frac{1}{2}\sum_{i}\nabla_{i}^{2}-\sum_{i}\frac{Z}{r_{i}}+\sum_{i<j}\frac{1}{|{\bf r}_{i}-{\bf r}_{j}|}$ (2)
is the standard electronic Hamiltonian for an atom with atomic number $Z$, and $V_{c}(r)$ is the potential for a hard-wall boundary at $r_{c}$
$V_{c}(r)=\begin{cases}0,\quad r<r_{c}\\ \infty,\quad r\geq r_{c}.\end{cases}$ (3)
As already mentioned in section 1, we carry out calculations on many-electron atoms within the context of DFT.26, 27 We expand the one-particle states also known as atomic orbitals (AOs) in a numerical basis set as
$\chi^{\sigma}_{nlm}(\mathbf{r})=R^{\sigma}_{nl}(r)Y_{lm}(\hat{\mathbf{r}}),$ (4)
where $Y_{\mathrm{lm}}(\hat{\mathbf{r}})$ are spherical harmonics; we will discuss the form of the radial expansion $R^{\sigma}_{nl}(r)$ below in section 3.
Following standard practice in numerical DFT calculations on atoms, we assume that the electron density is spherically symmetric, $n_{\sigma}({\bf r})=n_{\sigma}(r)$ for each spin $\sigma$. This simplification leads to an efficient algorithm, since the problem of solving the Kohn–Sham equations,27 i.e., the optimization of the $s$, $p$, $d$, and $f$ orbitals becomes a set of coupled one-dimensional problems.28, 29
Furthermore, we will assume integer occupations on the $s$, $p$, $d$ and $f$ shells in the aim to extract chemical understanding from the calculations. These integer occupations are then averaged over all magnetic sublevels to produce spherically symmetric total densities; this leads to fractional occupations on the individual spatial orbitals. The electron density is thus given simply as
$n_{\sigma}(r)=\sum_{nl}f^{\sigma}_{nl}[R^{\sigma}_{nl}(r)]^{2},$ (5)
where $f^{\sigma}_{nl}$ is the number of spin-$\sigma$ electrons on the $nl$ shell. We refer to refs. 28 and 29 for further discussion.
The above choice of theory has both a physical and practical rationale. The fractional occupation formalism automatically guarantees spherical degeneracy,28, 29 since all magnetic sublevels are assumed to behave identically. In contrast, DFT calculations with integer occupations of the magnetic sublevels often break spatial symmetry and fail to reproduce the degeneracy of atomic multiplets.28 Although several specialized approaches to reproduce atomic multiplets within a DFT style description have been proposed,30, 31, 32, 33, 34 they appear to be quite involved and have not gained wide adoption. We also wish to direct the reader to the discussion of Baerends et al.35 on atomic reference energies for DFT calculations.
3 Computational Details
FEM offers a straightforward way to the numerical solution of the radial functions $R^{\sigma}_{nl}$;36 a review of the employed FEM approach can be found in refs. 37 and 38. In short, the radial domain is first divided into $N_{\mathrm{elem}}$ segments $r\in[r_{i}^{\mathrm{start}},r_{i}^{\mathrm{end}}]$ called elements. An exponential radial grid $\{r_{i}\}_{i=0}^{N_{\mathrm{elem}}}$ is employed, so that the size of the elements increases with the distance from the nucleus.37, 38 A piecewise polynomial basis of shape functions $B_{n}(r)$ is then built up in each element.37, 38 Finally, the numerical radial basis functions to be used in eq. 4 are built from FEM shape functions $B_{n}(r)$ as
$R_{n}(r)=r^{-1}B_{n}(r),$ (6)
and the same radial basis set is used for all angular momenta $l$. This approach is robust and allows maximal flexibility: as the basis functions within one element have zero overlap to basis functions in other elements, the spatial representation can be adaptively refined, if necessary.
The endpoint of the last element is called the practical infinity, $r_{\infty}$, and all basis functions are built to vanish at $r_{\infty}$.37, 38 The physical interpretation of $r_{\infty}$ is that there is a hard-wall potential at this point; eq. 3 is therefore already built-in in the FEM approach, and this feature has been previously used in many studies, such as refs. 17 and 38. Thus, $r_{\infty}$ represents a physical parameter that expresses the location of the hard-wall boundary. In contrast, $r_{\infty}$ is a discretization parameter in studies of unconfined atoms (which represent the typical applications of FEM in the literature) that needs to be converged such that the obtained solution does not change if an even larger value is employed for $r_{\infty}$.
All calculations in this work were carried out with the free and open-source39 HelFEM program; the present implementation is publicly available on GitHub.40 Unless otherwise specified, the calculations employed shape functions defined by 15-node Lagrange interpolating polynomials (LIPs) specified by Gauss–Lobatto quadrature nodes; this corresponds to the use of a 14th order polynomial basis set. All calculations were converged to the complete basis set (CBS) limit with respect to the number of radial elements: radial elements were added until a $1\mu E_{h}$ precision in the total energy was achieved. We can thus be assured that we have found the numerically exact wave function.
We employ the optimal FEM implementation of the fractional occupation formalism recently described in refs. 28 and 29. All calculations employed the widely-used PBE exchange-correlation functional22, 23 as implemented in Libxc.41
We initiated the work by identifying the three lowest lying configurations of all elements with atomic number $1\leq Z\leq 54$ in the hard-wall potential for each $r_{c}\in\{1.0,1.2,1.4,\dots,10.0\}a_{0}$ with both spin-restricted and spin-polarized densities respectively, with a fixed angular momentum cutoff $l_{\mathrm{max}}=3$. The configurations were found with the automated algorithm introduced in ref. 28. The configurations considered for the spin-(un)restricted analysis were then chosen as the union of the low-lying configurations obtained from the spin-(un)restricted search over all of the above values of $r_{c}$.
We checked whether $g$ orbitals would become occupied for heavy atoms by repeating the configuration search for the Rb and Xe atoms with $l_{\mathrm{max}}=4$. However, no configurations with occupied $g$ orbitals were obtained in this search.
Thus being convinced to have found the low-lying configurations for the H–Xe atoms, we performed calculations on all these configurations with $r_{c}\in\{1.0,1.1,\dots,10.0\}a_{0}$. The wide range of studied values of $r_{c}$ allows us to observe many interesting evolutions. The states with $r_{c}=10.0a_{0}$ are similar to those of the unconfined atom, while $r_{c}=1.0a_{0}$ represents extreme confinement where even core orbitals experience confinement effects and the valence electrons are bound only by the confinement potential.
4 Results
4.1 Ionization energies of unconfined atoms
As a first step to validate our methods we compute ionization energies for the unconfined H–Xe atoms in the spin-restricted as well as the spin-polarized formalisms. The ionization energies can be calculated in two ways. First, the ionization energy is determined in the $\Delta$SCF method as the difference between the total energies of the neutral atom and its cation
$E_{+}=E_{\mathrm{atom}}-E_{\mathrm{ion}}.$ (7)
Second, while there is no Koopmans’ theorem42 for DFT as in Hartree–Fock, Janak’s theorem43 provides a justification for an analysis based on orbital energies.The ionization energy is then determined as
$E_{+}=\epsilon^{\mathrm{HOAO}},$ (8)
where $\epsilon^{\mathrm{HOAO}}$ is the eigenvalue of the highest occupied atomic orbital (HOAO).
(a) Ionization energies of the H–Xe atoms obtained with spin-restricted densities.
(b) Ionization energies of the H–Xe atoms obtained with spin-polarized densities.
Figure 1: Comparison of the ionization energies of gas-phase atoms computed with spin-restricted (fig. 1(a)) and spin-polarized (fig. 1(b)) densities in this work (TW) with $\Delta$SCF or Janak’s theorem against experimental values of Kramida et al.44.
The comparison of our computed values to the experimental values of Kramida et al.44 is depicted in fig. 1. The values obtained from Janak’s theorem are consistently and significantly lower than the experimental values and we will therefore not discuss them further. The $\Delta$SCF formalism produces ionization energies in good agreement with experiment. As expected, the results of the spin-polarized calculations are in better agreement with experiment than those from the spin-restricted calculations. Notable differences can be observed already for the H and He atoms. Spin-polarized calculations also faithfully reproduce the dip in the ionization energy beyond half filling of the $p$ shell, while spin-restricted calculations predict that the ionization energy increases monotonically with the filling of the $p$ shell.
4.2 Ionization radii
Having validated that the employed DFT method produces reliable ionization energies, we can proceed to computing ionization radii. The ionization radius $r_{+}$ is the value of $r_{c}$ for which it becomes energetically favorable for an atom to cast away its outermost electron. The analysis of Boeyens3 was based on non-relativistic Hartree–Fock calculations for $1\leq Z\leq 102$, and $r_{+}$ was identified via Koopmans’ theorem42 by finding the zero of the orbital energy of the outermost valence electron shell. However, as discussed in section 1, Boeyens3 did not employ a proper hard-wall potential, and did not consider the changing nature of the ground state configuration as a function of confinement.
4.2.1 $\Delta$SCF
The ionization radius $r_{+}$ for atom X is obtained within the $\Delta$SCF procedure by solving the nonlinear equation
$E({\rm X}^{+},r_{+})=E({\rm X},r_{+}),$ (9)
[…]
The second method is to realize that confinement can affect the ground state configurations of the atom and and its cation, and to compute the energies $E(X,r_{c})$ and $E(X^{+},r_{c})$ for the lowest configurations at each $r_{c}$. The resulting ionization radii are shown in table 2.
H 1.95 1.86 He 1.35 1.40
Li 4.17 4.06 Be 3.14 3.21 B 3.06 3.01 C 2.67 2.58 N 2.38 2.28 O 2.15 2.27 F 1.97 2.02 Ne 1.82 1.83
Na 4.39 4.31 Mg 3.71 3.77 Al 4.14 4.09 Si 3.65 3.55 P 3.28 3.18 S 2.99 3.12 Cl 2.76 2.81 Ar 2.56 2.57
K 5.26 5.18 Ca 4.61 4.66 Sc 4.37 4.46 Ti 4.54 4.27 V 4.43 4.09 Cr 4.08 3.94 Mn 3.95 3.96 Fe 3.83 3.79 Co 3.73 3.65 Ni 3.63 3.58 Cu 3.54 3.51 Zn 3.31 3.34 Ga 4.07 4.02 Ge 3.72 3.63 As 3.45 3.34 Se 3.22 3.34 Br 3.03 3.08 Kr 2.86 2.87
Rb 5.57 5.50 Sr 4.98 5.03 Y 4.72 4.82 Zr 4.88 4.55 Nb 4.52 4.38 Mo 4.37 4.23 Tc 3.59 4.13 Ru 3.40 4.04 Rh 3.24 3.29 Pd 3.09 3.11 Ag 3.88 3.85 Cd 3.66 3.69 In 4.42 4.37 Sn 4.09 4.01 Sb 3.84 3.74 Te 3.62 3.74 I 3.44 3.49 Xe 3.28 3.29
Table 1: Ionization radii $R$ in $a_{0}$ for the H–Xe atoms from spin-restricted (in red) and spin-polarized (in blue) non-relativistic $\Delta$SCF calculations with fractional occupations with configurations fixed to the ground states of the unconfined neutral atom and its cation.
H 1.95 1.86 He 1.35 1.40
Li 4.17 4.06 Be 3.14 3.21 B 3.06 3.01 C 2.67 2.58 N 2.38 2.28 O 2.15 2.27 F 1.97 2.02 Ne 1.82 1.83
Na 4.39 4.31 Mg 3.71 3.77 Al 4.14 4.09 Si 3.65 3.55 P 3.28 3.18 S 2.99 3.12 Cl 2.76 2.81 Ar 2.56 2.57
K 5.26 5.18 Ca 4.82 4.78 Sc 4.62 4.49 Ti 4.41 4.14 V 4.14 3.82 Cr 3.89 3.94 Mn 3.68 3.83 Fe 3.50 3.72 Co 3.36 3.46 Ni 3.23 3.26 Cu 3.54 3.51 Zn 3.31 3.34 Ga 4.07 4.02 Ge 3.72 3.63 As 3.45 3.34 Se 3.22 3.34 Br 3.03 3.08 Kr 2.86 2.87
Rb 5.57 5.50 Sr 5.13 5.07 Y 4.86 4.59 Zr 4.41 4.16 Nb 4.07 3.84 Mo 3.81 4.17 Tc 3.59 3.79 Ru 3.40 3.51 Rh 3.24 3.29 Pd 3.09 3.11 Ag 3.88 3.85 Cd 3.66 3.69 In 4.42 4.37 Sn 4.09 4.01 Sb 3.84 3.74 Te 3.62 3.74 I 3.44 3.49 Xe 3.28 3.29
Table 2: Ionization radii $R$ in $a_{0}$ for the H–Xe atoms from spin-restricted (in red) and spin-polarized (in blue) non-relativistic $\Delta$SCF calculations with fractional occupations, considering the $R$ dependence of the ground state of the neutral atom and its cation.
The ionization radius for the hydrogen atom is an interesting place to start the analysis. After all, one would expect the largest error to be made for this atom, as its exact ground state is fully spin polarized, while our calculation is for an atom with $1/2$ spin-up and $1/2$ spin-down electrons. The corresponding exact model is also analytically solvable: Sommerfeld and Welker45 obtained the ionization radius $1.8352a_{0}$ in their pioneering study. It is quite reassuring that our spin-polarized DFT value of $1.86a_{0}$ deviates from this exact value for H by only 1.4%, while the spin-restricted value of $1.95a_{0}$ exhibits a much larger deviation of 6.3%—which is still remarkably accurate.
The comparison of the results of the first and second methods proves interesting. Although inclusion of the relaxation effects does not appear to be important for most of the studied elements, large differences are observed in the pre-$d$ and $d$ blocks of the periodic table (Ca–Ni and Sr–Mo).
[…]
However, significant changes in energy upon relaxing the configuration may also occur for the cation, in which case the ionization radius can increase. Interestingly, such increases when the ground states are relaxed are observed for the Ca, Sc and Sr atoms in both the spin-restricted and spin-polarized cases, as well as the Y atom in the spin-restricted case. The corresponding cations exhibit a ground state crossing that moves the cross-over point of eq. 9 down in energy, and thereby the ionization radius to the right; see fig. 2 for an illustration of these four exceptional cases.
(a) Ca
(b) Sc
(c) Sr
(d) Y
Figure 2: Ground state crossings for the spin-polarized Ca, Sc and Sr atoms, and the spin-restricted Y atom in the hard-wall potential. Configurations for the neutral atoms shown as solid lines, and those corresponding to the monocation as dashed lines.
The differences observed in ionization radii with and without relaxing the electronic configuration underline the importance of considering various possible configurations as a function of confinement. Many transition metal atoms and their monocations exhibit ground state crossings already at weak confinement, due to the low-lying $s$-$d$ excitations in these systems (see section 4.3).
The spin-polarized and spin-restricted ionization radii are different for all atoms, which is not surprising given that either the neutral atom or its cation is open-shell. Interestingly, the differences between these two approaches are the smallest for the noble gas atoms, while alkali metals show much larger differences between spin-polarized and spin-unpolarized calculations. As the spin-polarized calculations also gave ionization energies in good agreement with experiment in section 4.1, the spin-polarized ionization radii ¸should be the most accurate.
4.2.2 Janak’s theorem
Another possibility is to employ a procedure similar to Boeyens3 who used Koopmans’ theorem to determine the ionization radii. As discussed in section 4.1, we can use Janak’s theorem, and the ionization radius $r_{+}$ is now found from
$\epsilon^{\text{HOAO}}({\rm X},r_{+})=0$ (10)
where $\epsilon^{\text{HOAO}}({\rm X})$ is the energy of the HOAO of the neutral atom X, with the hard wall boundary at $r_{c}=r_{+}$.
We report ionization radii obtained from calculations using Janak’s theorem without the TSA as tables in the SI, as we already found above in section 4.1 that these calculations failed to reproduce ionization energies in a reliable manner.
We note in passing that an exact solution to eq. 10 is problematic in some cases that exhibit a ground state crossing in the relevant range for $r_{+}$. A case in point is the V atom: the value of $r_{+}$ corresponds to the location of a ground state crossing between a state with $\epsilon^{\mathrm{HOAO}}(r)<0$ and a state with $\epsilon^{\mathrm{HOAO}}(r)>0$, which means that eq. 10 is not truly satisfied no matter how tightly $r_{+}$ is converged.
The origin of this problem is the present use of electron configurations with integer occupations. However, it has long been known that the use of non-integral occupation numbers for the $4s$ and $3d$ shells can result in a lower total energy for transition metal atoms in average-of-configurations Hartree–Fock calculations.46 This phenomenon is also topical for the present calculations with the PBE functional,47, 28 and has also been found in DFT calculations that do not employ any exchange-correlation functional.48 Calculations employing variationally optimized fractional occupation numbers would allow finding an exact solution to eq. 10. They are also expected to lead to somewhat different solutions of eq. 9. However, we do not consider such fractional occupations in this work, since integer occupation numbers are more commonly used and allow a clearer physical interpretation of the results.
Sen et al.13 computed atomic ionization radii with Janak’s theorem for the He–Ne atoms with exchange-only calculations based on Becke’s 1988 generalized-gradient functional.49 In contrast to our work, Sen et al.13 employed the transition state approximation (TSA) of Slater et al.46 in their calculations; the orbital energies were thus determined in calculations where half an electron was removed from the highest-lying orbital. Sen et al.13 compared this method to $\Delta$SCF calculations and found that the two methods yield ionization radii in close agreement to each other. Despite the differences between our approaches, we will see below that our results are in good agreement with those of Sen et al.13.
4.2.3 Comparison to literature radii
Numerically reliable ionization radii have been previously published by Boeyens3, Sen et al.13, and Garza et al.14. We compare our ionization radii from calculations based on the $\Delta$SCF method (section 4.2.1) and Janak’s theorem (section 4.2.2) with relaxed configurations against the earlier calculations in fig. 3.
We first note that the values of Boeyens3 are systematically—and considerably—larger than the results of any of the other calculations for $Z\geq 6$. These surprisingly large differences are likely explained by the lack of electron correlation in Hartree–Fock theory, which is known to lead to overestimation of atoms’ size compared to high-level wave function calculations, with DFT yielding close agreement with accurate theoretical reference values.50 The closeness of DFT densities with multideterminantal densities has also been noted on previously by Ortiz-Henarejos and San-Fabián51, for example.
Our calculations based on Janak’s theorem without the TSA systematically predict larger ionization radii than our $\Delta$SCF calculations, but they are still much smaller than the Hartree–Fock values of Boeyens3. Our $\Delta$SCF calculations are in excellent agreement with the limited data reported by Sen et al.13 in the spin-restricted formalism, who also found good agreement between their TSA method and $\Delta$SCF calculations in their work. Interestingly, our spin-polarized data appears to be in better agreement with the calculations of Sen et al.13 than our spin-restricted calculations. As was discussed in section 4.2.2, Sen et al.13 did not include correlation in their calculations, which likely explains the differences between our data.
Our $\Delta$SCF calculations are also in good agreement for the 9 atoms studied by Garza et al.14 with (apparently $\Delta$SCF) calculations based on the spin-unrestricted local-density approximation.
Since our $\Delta$SCF calculations yield ionization radii in good agreement with previous literature and ionization potentials of unconfined atoms in good agreement with experimental values (section 4.1), we are confident that our calculations offer a reliable basis for further studies of confined atoms with elaborate wave function methods. We further note that the dip in energy for the group 16 elements, which was found in the ionization energies in section 4.1, is also visible in the ionization radii in fig. 3.
(a) Spin-restricted densities
(b) Spin-polarized densities
Figure 3: Comparison of ionization radii computed in this work via $\Delta$SCF (blue circles) or Janak’s theorem (magenta squares) against the Hartree–Fock values of Boeyens3 (red triangles) and DFT values of Sen et al.13 (forest green diamonds) and Garza et al.14 (navy crosses). The calculations of this work were carried out either with spin-restricted (fig. 3(a)) densities, or spin-polarized (fig. 3(b)) densities.
4.3 Observed changes in electronic structure
Finally, we examine the behavior of the low lying states of the H–Xe atoms as a function of the hard-wall confinement radius $r_{c}\in[a_{0},10a_{0}]$. We pay special consideration to ground state crossings, which signify states that may be especially relevant for chemical bonding.
Because the various configurations’ energies span several orders of magnitude under the studied range of confinement radii, our graphical representation of the results rests on comparing the energy of the confined atom with electron configuration $i$ to that of the ground state configuration of the unconfined atom
$\Delta E_{i}(r_{0})=E_{i}(r_{0})-E_{0}^{\mathrm{unconfined}},$ (11)
where $i$ is the state in question. The plots are included in the SI, as we restrict the discussion in the main text to the observed qualitative changes in atomic electronic structure under confinement.
In addition to energies, we also report estimated atomic radii for various electronic configurations of the unconfined atoms. First, the location of the density maximum of the outermost orbital is a long-established estimator for the size of covalent atoms52
$r_{\mathrm{max}}=\mathrm{max}_{i}\left(\mathrm{argmax}\left[r^{2}n_{i}(r)\right]\right),$ (12)
where $n_{i}(r)$ is the $i$th orbital density. Second, since the electron density decays in the far valence region, a density threshold can be used to estimate the atom’s van der Waals size53
$n(r_{\mathrm{vdW}})=\epsilon.$ (13)
Bader et al.53 employed the threshold 0.002 electrons/bohr3, while Boyd54 showed that the criterion 0.001 electrons/bohr3 reproduces the same relative radii of atoms as those obtained with an even smaller value for the threshold. Rahm et al. also employed the threshold 0.001 electrons/bohr3 in their study of atomic and ionic radii55, 56 and we chose to adopt this value of the threshold, $\epsilon=10^{-3}$, as well.
We discovered in our analysis that $r_{\mathrm{vdW}}$ comes out surprisingly small for some diffuse states, such as the $1s^{1}3d^{1}$ configuration of He. The $r_{\mathrm{max}}$ estimator predicts this state to be considerably more extended than the $1s^{1}1p^{1}$ or $1s^{2}$ configurations ($r_{\mathrm{max}}=9.12a_{0}$, $3.68a_{0}$, and $0.56a_{0}$, respectively). Yet, $r_{\mathrm{vdW}}$ for the $1s^{1}3d^{1}$ configuration is smaller ($r_{\mathrm{vdW}}=2.08a_{0}$) than that for the $1s^{1}2p^{1}$ ($r_{\mathrm{vdW}}=4.10a_{0}$) or $1s^{2}$ ($r_{\mathrm{vdW}}=2.60a_{0}$) configurations.
To remedy this behavior, we introduce a third metric for atomic size by including the factor $r^{2}$ of the volume element in eq. 13 as
$r^{2}n(r_{\mathrm{r}^{2}\mathrm{vdW}})=\epsilon.$ (14)
This definition is somewhat more physical, since a spherical shell farther away from the nucleus can contain many more electrons, and thus having a small $n(r)$ does not alone guarantee that most of the electron density of the atom has been enclosed within $r_{\mathrm{r}^{2}\mathrm{vdW}}$. Due to the inclusion of the $r^{2}$ factor, eq. 14 appears to produce the largest estimates for atomic radii.
We note here that the radial density of electrons (the number of electrons found at $[r,r+{\rm d}r]$) is associated with an additional factor of $4\pi$. While the orbital density maximum in eq. 12 does not rely on a threshold and leads to the same radii whether or not this factor is included in the definition of $n(r)$, the threshold $\epsilon$ in eqs. 13 and 14 does depend on whether or not this factor has been included. Our implementation does not include the factor $4\pi$ in $n(r)$.
We provide a thorough analysis of the progression of the low lying states in hard-wall confinement in the SI. The analysis in the SI includes the three estimates of atomic radius for each configuration of the unconfined atom ($\epsilon=0.001$ for both $r_{\mathrm{vdW}}$ and $r_{\mathrm{r}^{2}\mathrm{vdW}}$), as well as their excitation energies relative to the ground state. We report the key observations from this analysis in the following.
H–He
We do not observe ground state crossings for the H or He atoms in the studied range of confinement radii. The $1s$ orbital remains occupied over the whole domain of studied confinement radii. Interestingly, both atoms have configurations with an occupied $3d$ orbital which are stable with respect to ionization in the unconfined atom (the configuration has a lower energy than that of the cation’s ground state), even though the $3d$ orbital is extremely diffuse with $\langle r\rangle=\mathcal{O}(10a_{0})$.
Li–O
The valence electrons of the Li–O atoms occupy the $2s$ orbital in the unconfined atom. Contrary to the H and He atoms, we observe ground state crossings where the $2s$ electrons hop to the $2p$ orbital as the confinement radius is decreased; for the Li atom this happens already at $r_{c}=3a_{0}$, but the location of the crossover point $r_{c}$ decreases in increasing charge and occurs for the O atom at $r_{c}=a_{0}$. We also note that the Li and O atoms have configurations with an occupied $3d$ orbital that are stable with respect to ionization in the unconfined atom, although the $3d$ orbital is again extremely diffuse with $\langle r\rangle=\mathcal{O}(10a_{0})$, similarly to H and He.
F
The valence electrons in the F atom occupy the $2s$ and $2p$ orbitals. One of the $2s$ electrons hops to the $2p$ orbital at $r_{c}=a_{0}$, which represents extreme confinement, thus fully filling the $2p$ orbital. Similarly to Li and O, F has a configuration with an occupied $3d$ orbital already in the unconfined atom, but the orbital is equally diffuse.
Ne
The Ne atom does not exhibit ground state crossings in the studied range of confinement radii. We note that also Ne has a configuration with an occupied $3d$ orbital which is stable with respect to ionization in the unconfined atom.
Na–Mg
The valence electrons of Na and Mg occupy the $3s$ orbital, which becomes unfavourable in confinement, like the $2s$ orbital in Li and Be. While in Li and Be the electrons hopped from $2s$ to $2p$, the $3s$ electrons of Na and Mg hop to the $3d$ orbital around $r_{c}\approx 2a_{0}$, instead, even though the $3d$ orbital has a larger extent than the $3p$ orbital in the unconfined atom. This suggests that the $3d$ orbital can be relevant for chemical bonding.
We observe that the [Ne]$4f$ configuration of the Na atom is stable with respect to ionization in the unconfined atom, and flips below the initial ground state in extreme confinement (close to $r_{c}=a_{0}$). The $4f$ orbital is highly diffuse in the unconfined atom ($\langle r\rangle=\mathcal{O}(20a_{0})$). Also the state [Ne]$3d^{1}4f^{1}$ of Mg flips below the initial ground state close to $r_{c}=a_{0}$.
Al–Ar
The valence electrons of the unconfined Al–Ar atoms occupy the $3s$ and $3p$ orbitals. Unlike the atoms in the previous period, we see that the $np$ orbital becomes unfavourable in confinement, gradually transferring all electrons to the $3d$ orbital when the confinement radius is decreased, the transition occurring between $r_{c}=2a_{0}$ and $r_{c}=1.5a_{0}$ (strong confinement). As the confinement radius is further decreased, both the $3s$ electrons follow. The behavior is thus different from that observed above for Li–Ne, likely because the $3d$ orbital is too high to be accessible in the lighter atoms.
K–Ca
Similarly to Na and Mg, the $ns$ valence electrons of the K and Ca atoms hop to the $3d$ orbital around $r_{c}\approx 4.5a_{0}$; the $3d$ orbital is thus already easily accessible, and likely relevant in chemistry. Further ground state crossings happen close to $r_{c}=a_{0}$ (extreme confinement), where the [Ar] core configuration is opened up and all the $3s$ and $3p$ electrons also move to the $3d$ orbital. We observe that both atoms have a configuration with an occupied $3d$ orbital that is bound relative to ionization in the gas-phase atom.
Sc–Ni
The transition metals feature strong competition between the $ns$ and $(n-1)d$ orbitals, which can be observed from the small excitation energies in the unconfined atom. The states [Ar]$3d^{n}$, $n\leq 10$ are stable with respect to ionization already in the unconfined atom, which explains why ground state crossings happen already in weak confinement.
The $4s$ electrons shift to the $3d$ orbital around $r_{c}=4a_{0}$ in the Sc–Ni atoms, as was also observed for K and Ca. The Sc and Ti atoms exhibit further ground state crossings in extreme confinement (between $r_{c}=a_{0}$ and $r_{c}=2a_{0}$), where some $3s$ and $3p$ electrons also transfer to the $3d$ orbital.
Cu–Zn
The unconfined Cu and Zn atoms already have a fully filled $3d$ orbital in their ground states, thus the $4s\rightarrow 3d$ shift observed earlier in $3d$ the transition metal series cannot happen for these atoms. The $4s$ electrons hop to the $4f$ orbital in extreme confinement (around $r_{c}=1.5a_{0}$), instead.
Ga–Kr
The ground state configurations of the unconfined Ga–Kr atoms have filled $4s$ and $3d$ orbitals and a variably filled $4p$ orbital. The $4s$ and $4p$ electrons hop to the $4f$ orbital as the confinement radius is decreased (between $r_{c}=1.5a_{0}$ and $r_{c}=a_{0}$), leading to a [Ar]$3d^{10}4f^{n}$ ground state in extreme confinement.
Rb–Sr
Like the alkali and earth alkali atoms in the previous period, the valence $ns$ electrons of the Rb and Sr atoms hop to the $(n-1)d$ orbital already around $r_{c}=5a_{0}$. However, unlike K and Ca, we also see a $4d\rightarrow 4f$ transfer when the confinement radius is further decreased to $r_{c}\approx 1.2a_{0}$, which represents extreme confinement. The $4s$ and $4p$ electrons also transfer to the $4f$ orbital below $r_{c}=1.2a_{0}$.
Y–Rh
The valence electrons of the unconfined Y–Rh atoms occupy the $ns$ and $(n-1)d$ orbitals in their ground states. The $5s$ valence orbital becomes unfavourable in confinement, and the electrons hop to the $4d$ orbital at confinement radii varying from $r_{c}=6.1a_{0}$ to $r_{c}=4.2a_{0}$, that is, already in weak confinement. Again, like for Cu–Sr, we observe electron transfers to the $4f$ orbital, and the $4d$ orbital is partially depopulated after the $5s$ orbital in confinement radii below $r_{c}=1.5a_{0}$, which already represents extreme confinement. Y and Zr transfer electrons from the $4p$ core orbital to the $4f$ orbital close to $r_{c}=a_{0}$.
Pd
The Pd atom has a particularly stable gas-phase ground state ([Kr]$4d^{10}$). We observe no ground state crossings for Pd in the studied range of confinement radii.
Ag–Cd
Ag and Cd already have a filled $4d$ orbital, while the $5s$ electrons hop to the $4f$ orbital around $r_{c}=2a_{0}$.
In–Xe
The In–Xe atoms have filled $5s$ and $4d$ orbitals and a $5p$ orbital with variable filling in the unconfined atom. The $5p$ electrons start to gradually shift to the $4f$ orbital around $r_{c}=2a_{0}$, after which both the $5s$ electrons follow analogously to Ga–Kr.
5 Summary and conclusion
To the best of our knowledge, a robust and systematic study of atoms in confinement has not been hitherto published in the literature. Such a study both needs to use a suitable numerical method and to consider the changing nature of the ground state as a function of the confinement, and we did not find a single publication that satisfies both these criteria, while considering an appreciable portion of the periodic table instead of just some specific subrows or groups.
We therefore took a fully numerical approach based on the finite element method (FEM) to study atoms in hard-wall confinement. We chose to represent the atoms with spherically averaged, spin-restricted or spin-polarized densities, and carried out calculations with the Perdew–Burke–Ernzerhof (PBE) exchange-correlation functional.22, 23 We compared the ionization energies of unconfined atoms predicted by spin-restricted or spin-polarized calculations to experiment, and observed the latter to be unsurprisingly more accurate, while even the spin-restricted calculations succeed in predicting periodic trends.
We then carried out a systematic study of the total energies of the ground and low lying excited states of the H–Xe atoms as well as their monocations as a function of the confinement radius $r_{c}$, paying special focus on how the ground state evolves as a function of confinement. We demonstrated the significance of considering confinement induced changes in the electronic structure by comparing ionization radii computed for fixed electronic configurations to those obtained in calculations, where the electronic configuration can relax as a function of the confinement radius. Depending on the atom, the relaxation can result in either an increase or a decrease of the ionization radius, and the largest differences are seen for the pre-$d$ and early $d$ elements.
We summarize our analysis of the calculations as follows.
  • Ground state crossings occur for the majority of the elements.
  • Valence $s$ electrons are highly unfavoured under strong confinement: already the Li and Be atoms show an $ns\to np$ electron transfer.
  • Shifts from the $ns$ and $np$ to the $3d$ orbital are observed for the elements of the second and third periods.
  • Shifts to the $4f$ orbital can be observed for Cu onwards in strong confinement.
Our systematic examination of the low lying configurations of the elements H–Xe highlights drastic changes in the electronic structure of confined atoms and provides a base for future studies with higher-accuracy wave function methods.
Like most studies on confined atoms, the present study was limited to the use of integer occupation numbers for the $s$, $p$, $d$, and $f$ shells to maximize physical interpretability. However, as we discussed in section 4.2.2, the use of non-integral occupation numbers could be useful for transition metals, for instance,46, 47, 28 and we hope to revisit such calculations with state-of-the-art optimization algorithms such as optimal damping57 in future work.
Our study was limited to the H–Xe atoms, because relativistic effects are well known to be essential for heavy elements.58, 59 Relativistic calculations of atoms in confinement should be feasible using the same approach, and such calculations might be carried out with the recent implementation of Čertík et al.60, for example.
{acknowledgement}
H.Å. thanks the Finnish Society for Sciences and Letters for funding. S.L. thanks the Academy of Finland for financial support under project numbers 350282 and 353749.
{suppinfo}
The SI PDF file contains the following data. Plots of the total energy of the low lying states for the H–Xe atoms and their monocations as a function of the confinement radius. Detailed analysis of the evolution of the ground state, and state crossings with the ground state configuration of the unconfined atom. Estimated atomic sizes and excitation energies relative to the ground state for the studied configurations in the unconfined atom. Values for ionization radii via Janak’s theorem for both the fixed and relaxed configuration approaches. All of the above sets of results are given both for the spin-restricted and spin-polarized formalisms.
The raw energies for all calculations are included in the comma separated values (CSV) format in a compressed archive.
{tocentry}[Uncaptioned image]